OpenStudy (jbedu):

I was working on some series problems, when I encountered this as part of "Mental Aptitude Test". The a(1) is a real number represents 'a' suffix '1'. Any help is appreciated. a(1) = 2 a(2) = 3 a(r+2) = [5/a(r)] + a(r+1) Find a(50) x a(51)

OpenStudy (karim728):

geometric series?

OpenStudy (jbedu):

Not sure if it is GP. I thought the product of 2 consecutive number should be in AP. But I may be wrong, as I could not proceed further.

OpenStudy (karim728):

i dont get this a(r+2) = [5/a(r)] + a(r+1)

OpenStudy (jbedu):

It means : \[a_{r+2} = \frac{ 5 }{ a _{r} } + a _{r+1}\]

OpenStudy (holsteremission):

Just out of curiosity, are you sure it's not \[a_{r+2}=\frac{5}{a_{r+1}}+a_r~?\]That appears to be much easier. Just making sure there's no typo.

OpenStudy (jbedu):

No its not, mate... If this is the case, its solvable ... it ends up as an AP with products of consecutive numbers. So the 50th term (\[a _{50} * a _{51}\]) equals to 251 as the initial number of the series will be 3 * 2 = 6. It is possible that its a typo in the book itself. I plan to write to the publisher. Thanks mate.